# MLAAD: The Multi-Language Audio Anti-Spoofing Dataset

### v3

@@ Abstract
Text-to-Speech (TTS) technology brings significant advantages, such as giving a voice to those with speech impairments, but also enables audio deepfakes and spoofs. The former mislead individuals and may propagate misinformation, while the latter undermine voice biometric security systems. AI-based detection can help to address these challenges by automatically differentiating between genuine and fabricated voice recordings. However, these models are only as good as their training data, which currently is severely limited due to an overwhelming concentration on English and Chinese audio in anti-spoofing databases, thus restricting its worldwide effectiveness.
In response, this paper presents the Multi-Language Audio Anti-Spoof Dataset (MLAAD), created using 54 TTS models, comprising 21 different architectures, to generate 163.9 hours of synthetic voice in 23 different languages. We train and evaluate three state-of-the-art deepfake detection models with MLAAD, and observe that MLAAD demonstrates superior performance over comparable datasets like InTheWild or FakeOrReal when used as a training resource. Furthermore, in comparison with the renowned ASVspoof 2019 dataset, MLAAD proves to be a complementary resource. In tests across eight datasets, MLAAD and ASVspoof 2019 alternately outperformed each other, both excelling on four datasets.
By publishing [^1] https://owncloud.fraunhofer.de/index.php/s/tL2Y1FKrWiX4ZtP MLAAD and making trained models accessible via an interactive webserver [^2] https://deepfake-total.com/, we aim to democratize anti-spoofing technology, making it accessible beyond the realm of specialists, thus contributing to global efforts against audio spoofing and deepfakes.
Index Terms:
text-to-speech, anti-spoofing, deepfake, voice, audio, biometrics
@@ I Introduction
To address these issues, we introduce our MLAAD dataset, a large-scale corpus of state-of-the-art audio fakes encompassing 23 languages. This dataset, based on the M-AILABS Speech Dataset [14], includes 163.9 hours of synthesized speech, created by 54 state-of-the-art TTS models. By training state-of-the-art deepfake detection models using MLAAD, we show that the resulting models achieve high performance scores in cross-dataset evaluations, indicating practicality in real-world conditions.
@@ III-C Statistics
The dataset encompasses 163.9 hours of synthesized speech across 23 languages, generated using 54 state-of-the-art TTS models, comprising 21 different architectures. We distribute only these synthesized samples, for the originals please refer to the M-AILABS dataset.
Figure 2 shows the model, dataset and language distribution within MLAAD, i.e., provides an overview of which TTS architectures, training datasets, and languages constitute the data. Here, “training datasets” indicates the data on which the TTS systems have been trained. These include both single speaker datasets [41, 45, 46, 64, 42], as well as the larger, multi-language and multi-speaker databases [43, 44, 14]. For some of the models, the training database is either unknown, non-existent (Griffin Lim), or composed of multiple datasets. The overall composition of TTS training datasets is highly diverse and includes both professional speakers [41] as well as audio recorded in community-driven initiatives [44].
@@ V Conclusion
In this paper, we present a new audio dataset, which comprises voice spoofs in 23 languages by 54 TTS models, composed of 21 architectures. When evaluating MLAAD’s usefulness as a training resource, we observe that MLAAD complements existing research. It consistently outperforms related work such as InTheWild or FakeOrReal with respect to the trained models’ cross-dataset generalization capability. Furthermore, in comparison with the renowned ASVspoof 2019 dataset, MLAAD proves to be a complementary resource. Across eight datasets, MLAAD and ASVspoof 2019 alternately outperformed each other, both excelling on four datasets: Training the SSL W2V2 model on ASVspoof19 results in the highest performance on ASVspoof21-DF, FakeOrReal, InTheWild, and MLAAD. Similarly, training SSL W2V2 or WhisperDF on MLAAD leads to optimal performance on ASVspoof19, ASVspoof21-LA, Voc.v, and WaveFake. We note that an evaluation of the multi-lingual anti-spoofing capabilities has to remain a topic for future work since there are no other comparable multi-lingual datasets against which to evaluate MLAAD. However, the strong results on WaveFake, which comprises both English and Japanese, suggest that the learning models profit from the diversity of languages in MLAAD.

### v4

@@ Abstract
Text-to-Speech (TTS) technology offers notable benefits, such as providing a voice for individuals with speech impairments, but it also facilitates the creation of audio deepfakes and spoofing attacks. AI-based detection methods can help mitigate these risks; however, the performance of such models is inherently dependent on the quality and diversity of their training data. Presently, the available datasets are heavily skewed towards English and Chinese audio, which limits the global applicability of these anti-spoofing systems.
To address this limitation, this paper presents the Multi-Language Audio Anti-Spoof Dataset (MLAAD), created using 59 TTS models, comprising 26 different architectures, to generate 175.0 hours of synthetic voice in 23 different languages.
We train and evaluate three state-of-the-art deepfake detection models with MLAAD and observe that it demonstrates superior performance over comparable datasets like InTheWild and FakeOrReal when used as a training resource. Moreover, compared to the renowned ASVspoof 2019 dataset, MLAAD proves to be a complementary resource. In tests across eight datasets, MLAAD and ASVspoof 2019 alternately outperformed each other, each excelling on four datasets.
By publishing [^1] https://deepfake-total.com/mlaad MLAAD and making a trained model accessible via an interactive webserver [^2] https://deepfake-total.com/ , we aim to democratize anti-spoofing technology, making it accessible beyond the realm of specialists, and contributing to global efforts against audio spoofing and deepfakes.
@@ I Introduction
To address these issues, we introduce our MLAAD dataset, a large-scale corpus of state-of-the-art audio fakes encompassing 23 languages. This dataset, based on the M-AILABS Speech Dataset [14], includes 175.0 hours of synthesized speech, created by 59 state-of-the-art TTS models. By training state-of-the-art deepfake detection models using MLAAD, we show that the resulting models achieve high performance scores in cross-dataset evaluations, indicating practicality in real-world conditions.
@@ III-C Statistics
The dataset encompasses 175.0 hours of synthesized speech across 23 languages, generated using 59 state-of-the-art TTS models, comprising 26 different architectures. We distribute only these synthesized samples, for the originals please refer to the M-AILABS dataset.
Figure 2, along with Table V and Table VI, presents the architecture, dataset, and language distribution within MLAAD, offering an overview of the TTS architectures, training datasets, and languages that constitute the dataset. Here, “training datasets” indicates the data on which the TTS systems have been trained. These include both single-speaker datasets [41, 45, 46, 64, 42], as well as larger, multi-language and multi-speaker databases [43, 44, 14]. For some of the models, the training database is either unknown, non-existent (e.g., Griffin Lim), or composed of multiple datasets. The overall composition of TTS training datasets is highly diverse and includes both professional speakers [41] and audio recorded through community-driven initiatives [44].
@@ V Conclusion
In this paper, we present a new audio dataset, which comprises voice spoofs in 23 languages by 59 TTS models, composed of 26 architectures. When evaluating MLAAD’s usefulness as a training resource, we observe that MLAAD complements existing research. It consistently outperforms related work such as InTheWild or FakeOrReal with respect to the trained models’ cross-dataset generalization capability. Furthermore, in comparison with the renowned ASVspoof 2019 dataset, MLAAD proves to be a complementary resource. Across eight datasets, MLAAD and ASVspoof 2019 alternately outperformed each other, both excelling on four datasets: Training the SSL W2V2 model on ASVspoof19 results in the highest performance on ASVspoof21-DF, FakeOrReal, InTheWild, and MLAAD. Similarly, training SSL W2V2 or WhisperDF on MLAAD leads to optimal performance on ASVspoof19, ASVspoof21-LA, Voc.v, and WaveFake. We note that an evaluation of the multi-lingual anti-spoofing capabilities has to remain a topic for future work since there are no other comparable multi-lingual datasets against which to evaluate MLAAD. However, the strong results on WaveFake, which comprises both English and Japanese, suggest that the learning models profit from the diversity of languages in MLAAD.